(4xy + 2x^2y) dx + (2x^2+3y^2) dy=0
1
Expert's answer
2020-10-13T17:18:12-0400

"\displaystyle(4xy + 2x^2) \mathrm{d}x + (2x^2+3y^2) \mathrm{d}y=0\\


M = (4xy + 2x^2), N = (2x^2+3y^2)\\


M_y = \frac{\partial M}{\partial y} = 4x\\

 
N_x = \frac{\partial N}{\partial x} = 4x\\


M_y = N_x, \\ \textsf{so the differential equation is exact.}\\


f(x, y) = \int 4xy + 2x^2\, \mathrm{d}x\\


f(x, y) = 2x^2y + \frac{2x^3}{3} + f(y) \\

\begin{aligned}
f(x, y) &= \int 2x^2+3y^2\, \mathrm{d}y
\\&= 2x^2y + y^3 + f(x)
\end{aligned}\\


\therefore f(x, y) = 2x^2y + \frac{2x^3}{3} + y^3 \\ \textsf{is the solution to the exact}\\\textsf{differential equation}"